OpenStudy (anonymous):

Zinc ore (zinc sulfide) is treated with sulfuric acid, leaving a solution with some undissolved bits of material and releasing hydrogen sulfide gas. If 10.8 g of zinc ore is treated with 50.0 mL of sulfuric acid (density 1.153 g/mL), 65.1 g of solution and undissolved material remains. In addition, hydrogen sulfide (density 1.393 g/L) is evolved. What is the volume (in liters) of this gas?

OpenStudy (abmon98):

(50.0 mL H2SO4) x (1.153 g H2SO4/mL) = 57.65 g H2SO4 total mass of reactants: 10.8 g + 57.65 g = 68.45 g By the Law of Conservation of Mass the total mass of products must also be 68.45 g. So: (68.45 g total) - (65.1 g leftovers) = 3.35 g H2S evolved (3.35 g H2S) / (1.393 g H2S/L) = 2.40 L H2S